# Chapter 1      Introduction to IoT

### 09 Dec 2023 v1

@@ 1.4.1 Generic Architecture of IoT
The three layer architecture is the generalized and most common architecture and several systems have integrated this architecture [64]. Although this multi layer architecture seems to be simple at first glance, but the functionalities of the network and application layer might get complex at times. For example, the network layer is not only responsible for data transmission but also provides data services such as data aggregation and processing etc. On the other hand, the application layer is not solely responsible for providing service to customers and users but also provides data analytic and conducts data mining etc. Therefore, in response to specific requirements, additional layers have been incorporated, building upon the fundamental layers. For instance, the four-layered or the five-layered architectures enhance the system’s flexibility. Nonetheless, the three-layer architecture serves as the foundation for all these variations. Furthermore, new-generation applications require shorter response times and low energy consumption as IoT devices have limited capacity [65, 66]. Therefore, researchers have utilized fog and cloud layers which are visualized in Fig. 1.4.
@@ 1.5 Key Technologies of IoT
To function properly, the Internet of Things requires a variety of technologies. Different technologies are deployed in various layers of an IoT architecture, and there are many different sorts of technologies, including hardware technology, software technology, and, most crucially, communication technology. This section discusses the many IoT technologies that are used to ensure the successful operation of an IoT system. Fig. 1.5 depicts the taxonomoy of IoT technologies.
@@ 1.5.1 Hardware Platforms
The major components of an IoT system are devices attached with sensors or wearable devices which are used to data collection. Therefore, various types of hardware platforms are used in order to build these sensor devices. Several key points need to be considered before selecting the hardware platforms such as what is the purpose of the IoT device, what kind of connectivity it requires and so on.
Two most accessible and popular hardware platforms are Raspberry Pi and Arduino. Raspberry Pi is a Linux based single board computer whereas Arduino makes different microcontrollers, open-source development tools, kits, and software. Both of them have strong data acquisition, processing ans storage capabilities and and provide both wireless and wired connectivity. However, in terms of power management Arduino is superior than Raspberry Pi because Raspberry Pi does not contain sleep or suspend modes for power utilization but Arduino does [67]. Intel has developed several development boards named Intel Galileo Gen and Intel Edison which uses Arduino IDE as development environment. Beagle Bone Black is a small sized Linux based board designed to build powerful IoT devices. Other than these platforms there are several other hardware platforms present in the markets such as Electric Imp 003, ARM mbed NXP LPC1768, WiPy and LoPy boards by Pycom company and many more [32].
@@ 1.5.2 Wireless Communication Technology
Given the vast number of devices in the current IoT network and the expectation of diverse connected devices in future IoT applications, there’s a urge to develop various technologies to facilitate their connectivity. This subsection explores the existing wireless technologies designed for IoT connectivity and categorize them into three groups.
Table 1.1: Comparison between popular short-range technologies based on various parameters. (Abbreviations: CSMA/CA (Carrier Sense Multiple Access with Collision Avoidance), CSMA/CD (Carrier Sense Multiple Access with Collision Detection), FHSS (Frequency Hopping Spread Spectrum), O-QPSK (Offset Quadrature Phase Shift Keying), OOK (On-Off Keying), PPM (Pulse Position Modulation), LBT (Listen Before Talk)
@@ 2.4 GHz
Bluetooth, ZigBee, LR-WPAN, and BS-ILC offer cost-effective communication solutions, whereas the overall cost of Wi-Fi and VLC may vary depending on usage. In terms of frequency bands, most technologies operate around 2.4 GHz, except for VLC, which utilizes a much larger frequency band. The frequency for BS-ILC varies by region. VLC and BS-ILC prioritize achieving high transmission rates of approximately 100 Gbps, while Wi-Fi transmission rates can vary across different standards. Notably, VLC and BS-ILC exhibit very low energy consumption, as indicated in the table.
Long Range Technologies
Long range technology in contrast with short range technologies refers to technologies that enable communication over extended distances. These technologies are designed to transmit data or signals over considerable ranges, typically much farther than short-range technologies like Bluetooth or Wi-Fi. Long-range technologies are commonly used in various applications such as long-distance wireless communication, remote monitoring, tracking, and more. They are essential for scenarios where data needs to be transmitted reliably over extensive geographical areas.
Long-range wireless technologies can be classified into two main categories: Mobile Communication Technology and LPWAN Technologies. Table 1.2 presents a detailed overview of various long-range technologies utilized in the IoT environment, organized based on several parameters such as frequency band, data rate, transmission range, energy consumption, cost, and more. it provides a summary of the technical specifications for various versions of Mobile Communication Technologies, which include 2G, 3G, 4G, 5G, and WiMax, as well as different versions of LPWAN Technologies, which encompass UNLICENSED LPWAN and LICENSED LPWAN.
It is evident from the table that when it comes to mobile communication technologies, 2G, 3G, and 4G operate within a frequency band ranging from approximately 850MHz to 2100MHz, while 5G utilizes a much broader frequency band of up to 80GHz. Additionally, all five technological variations, 5G offers the highest data rate, exceeding 20Gbps. In terms of energy efficiency, both 2G and 5G have low consumption rates, whereas WiMax exhibits high energy consumption. However, it’s worth noting that 3G, 4G, and 5G all come with higher implementation costs. In the UNLICENSED LPWAN category, the top contenders are LoRa and Sigfox, while the most promising options in the LICENSED LPWAN technology are LTE-M and NB-IoT (Narrowband Internet of Things). The key distinction highlighted in the table is that UNLICENSED LPWAN technology operates within unlicensed spectrum resources, leading to lower deployment costs, whereas LICENSED LPWAN technology relies on licensed spectrum resources, resulting in relatively higher expenses for both devices and deployment [68].
Emerging For Massive Connectivity
While current wireless IoT technologies have achieved some success in supporting various IoT applications, they still face challenges in meeting the future demands of IoT, where there will be a need to connect hundreds of billions of objects or things. For example, handling the connectivity among vast number of IoT devices with limited transmission payload and several random access protocols used by existing technologies often result in significant problems such as frequent access collisions, increased delays, and a high amount of signaling overhead for IoT devices. Furthermore, the constrained availability of wireless resources for connecting IoT devices creates shortages and inefficient utilization of these resources. Hence, there have been numerous ongoing initiatives to tackle the limitations of current technologies. Among these, the solutions CS, NOMA, mMIMO, and ML stand out as the most promising [68].
These technologies all have the potential to be incorporated into future IoT connectivity standards. However, it is essential to acknowledge that there are existing issues and constraints that must be resolved for their effective implementation. These emerging technologies offer the capability not only to support massive connectivity but also to deliver high reliability and low-latency transmissions. In the future, it is anticipated that the development of even more advanced technologies aimed at addressing critical IoT challenges. Simultaneously, efforts should focus on intelligently integrating existing and emerging technologies to unlock their full potential and optimize system performance [68].
Table 1.2: Comparison between popular long-range technologies based on various parameters. (Abbreviations: GSM (Global System for Mobile Communications), UMTS (Universal Mobile Telecommunications System), LTE-A (Long-Term Evolution Advanced), LTE-M (Long-Term Evolution for Machines), 3GPP (3rd Generation Partnership Project), TDMA (Time Division Multiple Access), CDMA (Code Division Multiple Access), FDMA ( Frequency Division Multiple Access), WCDMA (Wideband Code Division Multiple Access), OFDMA (Orthogonal Frequency Division Multiple Access), SC-FDMA (Single Carrier Frequency Division Multiple Access), CP-OFDM (Cyclic Prefix Orthogonal Frequency Division Multiplexing), TDD (Time Division Duplex), FDD (Frequency Division Duplex), ALHOA (Adaptive Link Hopping On Air), QPSK (Quadrature Phase Shift Keying), BPSK (Binary Phase Shift Keying), QAM (Quadrature Amplitude Modulation), GFSK (Gaussian Frequency Shift Keying), DBPSK (Differential Binary Phase Shift Keying), CSS (Chirp Spread Spectrum), OFDM (Orthogonal Frequency Division Multiplexing), GMSK (Gaussian Minimum Shift Keying))
@@
Moderae to high Moderae to high
@@ 1.5.3 Cloud Solutions
IoT cloud solutions offer various services like real-time data collection, data transmission, monitoring, data analytics, improved decision-making, and device management. These services are available on a pay-as-you-use basis, allowing users to pay only for the services they actually use. Cloud services are becoming increasingly popular, and cloud-based architectures are discussed in detail in Section 1.4. Cloud platforms can be integrated into numerous domains, including healthcare, smart cities, agriculture, education, and supply chain management. For simplicity purpose, Table 1.3 presents several popular platforms in the agriculture domain [32].
@@ 1.5.4 Hardware and Software Technology
This category encompasses a range of hardware and software technologies responsible for different aspects of IoT communication channels. Several technologies, such as NFC and M2M, are considered as both hardware and software technologies, whereas RFID is classified as a hardware-only technology. The subsequent sections will provide brief insights into these technologies.
Hardware Technology
@@
RFID, or Radio-Frequency Identification which has also been introduced in sub section 1.2.2, is a technology that comprises of one or more readers and several RFID tags. These tags are small microchips, have unique codes and can be on items like products in a store or even access cards. When a device called an RFID reader sends out electromagnetic radio waves, the tags respond with their codes. The reader captures these codes and sends them to a computer, which can then figure out what the tags belong to. This technology is used in various ways, like keeping track of inventory in stores, allowing access to secure areas, or monitoring packages as they move through the delivery process. Hitachi has developed a tag with dimensions that comes in three configurations Passive Reader Active Tag (PRAT), Active Reader Passive Tags (ARPT) and Active Reader Active Tag (ARAT). An Electronic Product Code (EPC) is a typical piece of information stored in an RFID tag. These codes make it possible to uniquely track objects because they are encoded on RFID tags [69].
@@
Near Field Communication (NFC) is a short range wireless connectivity technology that uses magnetic field induction to establish communicating between devices when they are brought within near proximity of each other without the need of prior connection establishment. It is an extension of RFID technology with additional and more dynamic features by integrating smartphones. Most modern phones are equipped with NFC chips and support applications like Apple Pay and Google Pay. NFC operates in the unlicensed Radio Frequency band at 13.56MHz. Its typical range is about 20 meters, with the actual distance often determined by the size of the antenna in the device. NFC technology is expected to play a vital role in the future of IoT by providing a wireless connection tool to link with other smart objects. For instance, using NFC on a mobile device, a user can turn their phone into various objects, like using it as a credit card for transactions [69].
Machine to Machine (M2M)
Machine-to-Machine (M2M) communication is gaining popularity and involves direct communication between computers, embedded processors, smart sensors, actuators, and mobile devices. It comprises four main components: sensing, heterogeneous access, information processing, and applications & processing. In practical terms, M2M functions within a five-part framework:
  1. 1.
    M2M Device: This is a device capable of responding to requests for data stored within it.
  2. 2.
    M2M Area Network (Device Domain): It provides the connectivity that links M2M Devices with M2M Gateways.
  3. 3.
    M2M Gateway: These gateways utilize M2M capabilities to ensure that M2M Devices can interact and connect to the communication network.
  4. 4.
    M2M Communication Networks (Network Domain): These networks facilitate communication between the M2M Gateway(s) and M2M applications.
  5. 5.
    M2M Applications: This layer serves as middleware, where data passes through various application services and is utilized by specific business processing engines.
Machine-to-Machine (M2M) technology finds applications across various sectors, including healthcare, smart robotics, cyber transportation systems (CTS), manufacturing systems, smart home technologies, and smart grids [69].
@@
Vehicle-to-Vehicle Communications (V2V) treats each vehicle as a node and enables wireless data exchange between vehicles regarding their speed, location, and more. This network allows vehicles to broadcast and receive messages in all directions. Vehicles equipped with appropriate software, often referred to as safety applications, can use messages from nearby vehicles to detect potential collision risks in real-time. There are two types of communication within this network: one is between vehicles, and the other involves road infrastructure. However, the structure or arrangement of this communication is flexible as vehicles move from one location to another. This network can be divided into four main categories: safety and collision avoidance, traffic infrastructure management, vehicle telematics, and entertainment services with Internet connectivity [69].
@@ 1.6.1 Smart Cities
The concept of smart cities can be seen as a complex IoT paradigm, where the management of public affairs incorporates the introduction of Information and Communication Technology (ICT) solutions [70]. A smart city utilizes public resources to digitize the city and enhance the quality of life. A real-world example of a smart city implementation is “Padova Smart City”, which has been put into practice in the city of Padova, Italy [71]. Barcelona and Stockholm are two noteworthy examples of smart cities. Barcelona has embarked on the CityOS project, with the primary goal of creating a centralized operating system to manage all the smart devices and services available within the city. Their focus has predominantly centered on enhancing smart transportation and water systems. Similarly, Stockholm has also placed significant emphasis on these two domains and holds the distinction of being one of the pioneering cities to implement the concept of congestion charges, wherein users are imposed fees for entering congested areas [4].
Smart cities are characterized by technology-based infrastructure and highly efficient smart public services. These smart cities are divided into multiple subsystems, and some of them are described below. In addition, Table 1.4 summarises related research works in this domain.
@@
Applications Related Works
Structural health of buildings [72, 70]
Waste management [70]
Air quality and noise monitoring [73, 74]
Smart Transport & Traffic congestion
@@
Smart grid [79, 80]
Smart water systems [81, 4]
@@
This service involves continuous monitoring of areas prone to various external factors, and measuring the health of any building. IoT sensors connected to the buildings can store information about the building’s strength, which will help analyze how sturdy the building is or if it requires any refinement [72]. Additionally, various kinds of sensors can be employed depending on the usage, such as sensors for tracking vibrations to measure the stress on the building, temperature and humidity sensors, or other types of atmospheric sensors to assess the level of pollution in an area [70]. Employing IoT in this area can reduce manual labor, where humans have to manually assess the building’s health and environmental conditions, and it can also reduce the overall cost to a great extent.
@@
Waste management is a crucial operation in any city, whether it is considered a smart city or not, as it directly impacts the livability of the area. Thus, an efficient waste management system is essential for any society. Integrating IoT into waste management offers multiple benefits. It enables the detection of waste levels and real-time tracking of garbage truck routes, leading to more efficient route planning. Moreover, it can streamline manual labor involved in waste separation and monitor the disposal process. These tasks are achieved using sensors on garbage vehicles connected to a central software system that analyzes and controls the system based on the data collected by the sensors [70]. This approach reduces the cost of manual processes and enhances recycling management.
@@
Creating a healthy and safe environment for all living beings is crucial, and monitoring air quality and noise levels in any area is a key part of achieving this goal. Various environmental sensors, including soil sensors, temperature and humidity sensors, and gas sensors, can detect the presence of toxic and pollutant substances in the air and assess pollution levels. This information enables local authorities to control pollution, implement effective measures to reduce pollutant levels, identify highly polluted or toxic areas, and designate suitable locations for outdoor activities with good air quality [73, 74]. Likewise, it is essential to maintain a balanced noise level for all living beings in society, including humans and animals. Noise sensors can be used to measure the decibel levels, and the data collected can help the central authority identify noisy areas and regulate noise levels to keep them within acceptable limits.
Smart Transport & Traffic Congestion
@@
In today’s world, traffic congestion is a widespread issue affecting nearly every country and city, particularly as more than half of the world’s population now resides in urban areas [4]. To tackle this issue, many cities have adopted IoT solutions to establish smart transportation systems aimed at managing traffic congestion.These initiatives include smart traffic lights and smart parking systems, which collectively enhance transportation capacity and improve safety and speed for travelers.
The primary advantages of smart transportation systems are reducing traffic congestion, ensuring hassle-free travel, and facilitating easy parking. Moreover, these systems enable quicker responses in case of accidents and contribute to accident reduction by effectively managing traffic flow [4]. A variety of sensor technologies are employed to achieve these objectives, such as accelerometers to measure speed, RFIDs for vehicle identification, GPS sensors for location tracking, gyroscopes for direction detection, and cameras for recording traffic patterns and vehicle movements. Practical implementations of these sensors are elaborated upon below.
  1. 1.
    Traffic Management and Monitoring Applications
    In a smart transportation system, vehicles are seamlessly interconnected via wireless networks. Each vehicle is equipped with essential IoT sensors, such as RFID and GPS. Traffic congestion is identified through the use of GPS and accelerometers, which are capable of tracking vehicle movements across different areas. The collected data is then analyzed by mapping systems to pinpoint congested areas. This information is subsequently transferred to the driver, providing guidance on selecting the optimal route and avoiding potential traffic jams [4, 75].
  2. 2.
    Application to Ensure Safety
    Numerous IoT applications have been developed with a focus on monitoring driver behavior continuously to enhance road safety. This is achieved by detecting signs of drowsiness or fatigue in drivers and providing appropriate recommendations, such as suggesting rest breaks. Various IoT applications utilize techniques like eye movement detection, facial expression analysis, and pressure sensing. These applications are typically integrated into smartphones, equipped with sensors like gyroscopes, accelerometers, and GPS, which analyze the collected data to make informed decisions. Additionally, through remote vehicle monitoring, it becomes feasible to track factors such as speed and tire pressure. This information allows centralized systems to make informed decisions aimed at preventing potential accidents [76].
  3. 3.
    Parking Guidance and Information System
    To address vehicle parking challenges, a range of IoT-based applications have been introduced. The primary goal of these applications is to assist drivers in locating available parking spaces remotely, eliminating the need to spend valuable time searching for vacant spots. This, in turn, contributes to reducing traffic congestion [77]. Parking guidance and information (PGI) applications employ barcodes to track the number of vehicles entering or leaving a parking area. Furthermore, parking zones are equipped with cameras and sensors that detect and monitor vehicles, facilitating the identification of vacant and suitable parking spots for incoming vehicles [78].
  4. 4.
    Smart Traffic Lights
    Smart traffic lights utilize data collected from sensors to make decisions based on the degree of traffic congestion. These intelligent systems can serve various purposes, such as prioritizing public transportation vehicles or implementing special controls for emergency service vehicles when needed [4].
  5. 5.
    Applications for Detecting Accidents
    These applications have the capability to identify unusual situations, such as road accidents, through the use of accelerometer sensors or the assistance of acoustic information. Additionally, they can transmit images of the surrounding area to the central authority. Based on this information, the authority can make informed decisions regarding the extent of medical assistance required.
@@
A traditional grid system is an electrical grid that includes transmission lines, transformers, and various communication utilities, responsible for delivering electricity from power plants to homes or businesses. One significant limitation of the traditional grid is its one-way communication, which prevents power plants from efficiently responding to increasing power demands from consumers. To address this challenge, the smart grid establishes a two-way communication system between utilities and consumers which enables more effective management of economic, sustainable, and secure power resources [79].
By integrating IoT into grid systems, houses and businesses can be equipped with smart meters that monitor energy generation, storage, and consumption, transmitting this data to the smart grid. The smart grid, which allows for two-way communication between utilities and consumers, helps consumers in better understanding the power consumption, reports real-time increases in energy demand and calculates electricity pricing for customers, departing from the traditional monthly billing cycle [80]. Additionally, with the aid of fog/edge computing, utility providers can efficiently dispatch energy from the grid, enabling consumers to reduce their energy consumption, utilize resources efficiently, and ultimately reduce their electricity costs.
@@
Water is one of the most critical natural resources, and its scarcity is a prevalent issue in many parts of the world. Therefore, implementing smart water systems is not a luxury but a necessity. The primary role of smart water systems is to monitor, measure, and efficiently distribute water usage. One notable model in this field is the smart water meter designed by Hauber-Davidson and Idris [81]. These meters can detect water inflow and outflow and identify any potential leaks. Smart water meters can also be used in conjunction with data from smart river sensors and weather information to aid in flood prediction [4].
1.6.2 Medical and Healthcare
The healthcare sector has experienced remarkable advancements through the integration of IoT, offering solutions to real-life healthcare challenges and enhancing people’s lifestyles. Researchers have proposed various applications in healthcare that utilize wearable sensor devices to monitor patients’ health, diagnose diseases, issue emergency alerts, and notify users when necessary. Remote monitoring saves time for both patients and doctors while reducing overall healthcare costs. Furthermore, sharing data collected from these wearable devices with healthcare researchers contributes to the development of safer and more timely healthcare solutions and aids in the discovery of cures and vaccines for emerging diseases. This section explores several IoT applications in healthcare and a summary of the associated research papers is presented in Table 1.5.
@@ 1.6.2 Medical and Healthcare
Applications Related Works
ECG Monitoring [82, 83]
Glucose level monitoring [84, 85]
Temperature monitoring [86, 87]
Blood pressure monitoring [88, 89]
Oxygen saturation monitoring [90, 91]
Mood monitoring [92, 93]
Medication management [94, 95]
Wheelchair management [96, 97]
Rehabilitation system [98, 99]
Fitness [100, 101]
Other notable applications

### 12 Jan 2024 v2

@@ 1.4 IoT Architecture
The three-layer architecture is the generalized and most common architecture, and several systems have integrated this architecture [64]. Although this multi-layer architecture seems simple at first glance, the functionalities of the network and application layers might get complex at times. For example, the network layer is not only responsible for data transmission but also provides data services such as data aggregation and processing, etc. On the other hand, the application layer is not solely responsible for providing service to customers and users but also provides data analysis, conducts data mining, etc. Therefore, in response to specific requirements, additional layers have been incorporated, building upon the fundamental layers. For instance, the four-layered or five-layered architectures enhance the system’s flexibility. Nonetheless, the three-layer architecture serves as the foundation for all these variations. Furthermore, new-generation applications require shorter response times and low energy consumption as IoT devices have limited capacity [65, 66]. Therefore, researchers have utilized fog and cloud layers, which are visualized in Fig. 1.4.
@@ 1.5 Key Technologies of IoT
IoT requires a variety of technologies to operate, which are deployed in various layers of an IoT architecture, and there are many different sorts of technologies, including hardware technology, software technology, and, most crucially, communication technology. This section discusses crucial IoT technologies that are used to ensure the successful operation of an IoT system. Fig. 1.5 depicts the taxonomy of IoT technologies.
@@ 1.5.1 Hardware Platforms
The major components of an IoT system are devices attached to sensors or wearable devices that are used for data collection. Therefore, various types of hardware platforms are used in order to build these sensor devices. Several key points need to be considered before selecting the hardware platforms, such as the purpose of the IoT device or the type of connectivity it requires.
The two most accessible and popular hardware platforms are the Raspberry Pi and Arduino. Both of them have strong data acquisition, processing, and storage capabilities and provide both wireless and wired connectivity. However, in terms of power management, Arduino is superior to Raspberry Pi because Raspberry Pi does not contain sleep or suspend modes for power utilization, while Arduino does [67]. Intel Galileo Gen and Intel Edison are examples of using an Arduino IDE.
@@ 1.5.2 Wireless Communication Technology
Given the vast number of devices in the current IoT network and the expectation of diverse connected devices in future IoT applications, there’s an urge to develop various technologies to facilitate their connectivity. This subsection explores the existing wireless technologies designed for IoT connectivity and categorizes them into three groups.
Table 1.2: Comparison between popular short-range technologies based on various parameters.
@@ 2.4 GHz
Table 1.3: Comparison between popular long-range technologies based on various parameters.
@@
Moderate to high Moderate to high
Bluetooth, ZigBee, LR-WPAN, and BS-ILC offer cost-effective communication solutions, whereas the overall cost of Wi-Fi and VLC may vary depending on usage. In terms of frequency bands, most technologies operate around 2.4 GHz, except for VLC, which utilizes a much larger frequency band. The frequency of BS-ILC varies by region. VLC and BS-ILC prioritize achieving high transmission rates of approximately 100 Gbps, while Wi-Fi transmission rates can vary across different standards. Notably, VLC and BS-ILC exhibit very low energy consumption, as indicated in the table.
Long Range Technologies
Long-range technologies are designed to transmit data or signals over considerable distances, typically much farther than short-range technologies like Bluetooth or Wi-Fi. Long-range technologies are commonly used in various applications, such as long-distance wireless communication, remote monitoring, tracking, and more, and are essential for scenarios where data needs to be transmitted reliably over extensive geographical areas. These technologies can be classified into two main categories: mobile communication technology and LPWAN technologies. Table 1.3 presents a detailed overview of various long-range technologies utilized in the IoT environment. It provides a summary of the technical specifications for various versions of mobile communication technologies as well as different versions of LPWAN technologies.
It is evident from the table that when it comes to mobile communication technologies, 2G, 3G, and 4G operate within a frequency band ranging from approximately 850MHz to 2100MHz, while 5G utilizes a much broader frequency band of up to 80GHz. Additionally, of all five technological variations, 5G offers the highest data rate, exceeding 20Gbps. In terms of energy efficiency, both 2G and 5G have low consumption rates, whereas WiMax exhibits high energy consumption. However, it is worth noting that 3G, 4G, and 5G all come with higher implementation costs. In the UNLICENSED LPWAN category, the top contenders are LoRa and Sigfox, while the most promising options in the LICENSED LPWAN technology are LTE-M (Long-Term Evolution for Machines) and NB-IoT (Narrowband IoT). The key distinction highlighted in the table is that UNLICENSED LPWAN technology operates within unlicensed spectrum resources, leading to lower deployment costs, whereas LICENSED LPWAN technology relies on licensed spectrum resources, resulting in relatively higher expenses for both devices and deployment [68].
Emerging For Massive Connectivity
While current wireless IoT technologies have achieved some success in supporting various IoT applications, they still face challenges in meeting the future demands of IoT. For example, handling the connectivity among a vast number of IoT devices with a limited transmission payload and several random access protocols used by existing technologies often results in significant problems such as frequent access collisions, increased delays, and a high amount of signaling overhead for IoT devices. Furthermore, the constrained availability of wireless resources for connecting IoT devices creates shortages and inefficient utilization of these resources. Hence, there have been numerous ongoing initiatives to tackle the limitations of current technologies. Among these, the solutions CS, NOMA, mMIMO, and ML stand out as the most promising [68].
These emerging technologies offer the capability not only to support massive connectivity but also to deliver high reliability and low-latency transmissions. However, it is essential to acknowledge that there are existing issues and constraints that must be resolved for their effective implementation. It is anticipated that the development of even more advanced technologies will address critical IoT challenges. Simultaneously, efforts should focus on intelligently integrating existing and emerging technologies to unlock their full potential and optimize system performance [68].
@@ 1.5.3 Cloud Solutions
IoT cloud solutions offer various services like real-time data collection, data transmission, monitoring, data analytics, improved decision-making, and device management. These services are available on a pay-as-you-use basis, allowing users to pay only for the services they actually use. Cloud platforms can be integrated into numerous domains, including healthcare, smart cities, agriculture, education, and supply chain management. For simplicity, Table 1.4 presents several popular platforms in the agriculture domain [32].
@@ 1.5.4 Hardware and Software Technology
This category encompasses a range of hardware and software technologies responsible for different aspects of IoT communication channels. Several technologies, such as NFC and M2M, are considered both hardware and software technologies, whereas RFID is classified as a hardware-only technology. The subsequent sections will provide brief insights into these technologies.
@@
RFID is a technology that consists of one or more readers and several RFID tags. These tags are small microchips, have unique codes, and can be on items like products in a store or even access cards. When an RFID reader device sends out electromagnetic radio waves, the tags respond with their unique Electronic Product Codes (EPCs) [69]. The reader captures these codes and sends them to a computer, which can then figure out what the tags belong to. This technology is used in various ways, like keeping track of inventory in stores, allowing access to secure areas, or monitoring packages as they move through the delivery process.
@@
NFC, an extension of RFID technology, is a short-range wireless connectivity technology that uses magnetic field induction to establish communication between devices when they are brought within close proximity of each other without the need for prior connection establishment. NFC chips are equipped in most modern phones, supporting applications like Apple Pay and Google Pay. NFC operates in the unlicensed radio frequency band at 13.56MHz. Its typical range is about 20 meters, with the actual distance often determined by the size of the antenna in the device. It is expected that NFC technology will play a vital role in the future of IoT by providing a wireless connection tool to link with other smart objects. For instance, using NFC on a mobile device, a user can transform their phone into various objects, such as using it as a credit card for transactions [69].
M2M
M2M communication is gaining popularity and involves direct communication between computers, embedded processors, smart sensors, actuators, and mobile devices. It comprises four main components: sensing, heterogeneous access, information processing, and applications and processing. In practical terms, M2M functions within a five-part framework: a device for responding to requests, gateways to interact and connect, an area network for providing connectivity between devices and gateways, applications serving as middleware, and a communication network to facilitate communication between gateways and applications. M2M technology is applied across various sectors, including healthcare, smart robotics, cyber transportation systems (CTS), manufacturing systems, smart home technologies, and smart grids [69].
@@
V2V Communications treats each vehicle as a node and enables wireless data exchange (omnidirectional) between vehicles regarding their speed, location, and more. Vehicles equipped with appropriate software, often referred to as safety applications, can use messages from nearby vehicles to detect potential collision risks in real-time. There are two types of communication within this network: one is between vehicles, and the other involves road infrastructure. However, the structure or arrangement of this communication is flexible as vehicles move from one location to another. This network can be divided into four main categories: safety and collision avoidance, traffic infrastructure management, vehicle telematics, and entertainment services with Internet connectivity [69].
@@ 1.6.1 Smart Cities
The concept of smart cities can be seen as a complex IoT paradigm where the management of public affairs incorporates the introduction of information and communication technology (ICT) solutions [70]. A smart city utilizes public resources to digitize the city and enhance the quality of life. A real-world example of a smart city implementation is “Padova Smart City”, which has been put into practice in the city of Padova, Italy [71]. Barcelona and Stockholm are two noteworthy examples of smart cities. Barcelona has embarked on the CityOS project, with the primary goal of creating a centralized operating system to manage all the smart devices and services available within the city. Their focus has predominantly been on enhancing smart transportation and water systems. Similarly, Stockholm has also placed significant emphasis on these two domains and holds the distinction of being one of the pioneering cities to implement the concept of congestion charges, wherein users are imposed fees for entering congested areas [4]. In addition, Table 1.5 summarizes related research works in this domain.
@@
Applications Related Works Applications Related Works
Structural health of buildings [72, 70] Waste management [70]
Air quality and noise
monitoring
[73, 74]
Smart Transport and
Traffic congestion
@@
Smart grid [79, 80] Smart water systems [81, 4]
@@
This service involves continuous monitoring of areas prone to various external factors and measuring the health of any building. IoT sensors connected to the buildings can store information about the building’s strength, which will help analyze how sturdy the building is or if it requires any refinement [72]. Depending on the usage, one can employ various kinds of sensors, including those for tracking vibrations to measure the stress on the building, temperature and humidity sensors, or other types of atmospheric sensors to assess the level of pollution in an area [70]. Employing IoT in this area can reduce manual labor, where humans have to manually assess the building’s health and environmental conditions, and it can also reduce the overall cost to a great extent.
@@
Waste management is a crucial operation in any city, whether it is considered a smart city or not, as it directly impacts the livability of the area. Thus, an efficient waste management system is essential for any society. Integrating the IoT into waste management offers multiple benefits. It enables the detection of waste levels and real-time tracking of garbage truck routes, leading to more efficient route planning. Moreover, it can streamline the manual labor involved in waste separation and monitor the disposal process. Sensors on garbage vehicles connected to a central software system achieve these tasks by analyzing and controlling the system based on the data collected [70]. This approach reduces the cost of manual processes and enhances recycling management.
@@
Creating a healthy and safe environment for all living beings is crucial, and monitoring air quality and noise levels in any area is a key part of achieving this goal. Various environmental sensors, including soil sensors, temperature and humidity sensors, and gas sensors, can detect the presence of toxic and pollutant substances in the air and assess pollution levels. This information enables local authorities to control pollution, implement effective measures to reduce pollutant levels, identify highly polluted or toxic areas, and designate suitable locations for outdoor activities with good air quality [73, 74]. Likewise, it is essential to maintain a balanced noise level for all living beings in society, including humans and animals. Using noise sensors to measure decibel levels, the central authority can collect data to identify noisy areas and regulate noise levels to keep them within acceptable limits.
Smart Transport and Traffic Congestion
@@
In today’s world, traffic congestion is a widespread issue affecting nearly every country and city, particularly as more than half of the world’s population now resides in urban areas [4]. To tackle this issue, many cities have adopted IoT solutions to establish smart transportation systems aimed at managing traffic congestion. These initiatives include smart traffic lights [4] and smart parking system [77, 78], which collectively enhance transportation capacity and improve safety and speed for travelers. The primary advantages of smart transportation systems are reducing traffic congestion, ensuring hassle-free travel, and facilitating easy parking. Moreover, these systems enable quicker responses in case of accidents and contribute to accident reduction by effectively managing traffic flow [4]. These objectives are achieved through the use of a variety of sensor technologies, including accelerometers for measuring speed, RFIDs for vehicle identification, GPS sensors for location tracking, gyroscopes for direction detection, and cameras for recording traffic patterns and vehicle movements. Some real-world uses of these sensors can be seen in applications for managing and monitoring traffic [4, 75], applications to ensure safety [76], and application for detecting accidents.
@@
A traditional grid system is an electrical grid that includes transmission lines, transformers, and various communication utilities responsible for delivering electricity from power plants to homes or businesses. One significant limitation of the traditional grid is its one-way communication, which prevents power plants from efficiently responding to increasing power demands from consumers. To address this challenge, the smart grid establishes a two-way communication system between utilities and consumers, which enables more effective management of economic, sustainable, and secure power resources [79]. Integrating IoT into grid systems enables equipping houses and businesses with smart meters that monitor energy generation, storage, and consumption, and transmit this data to the smart grid.
@@
Water is one of the most critical natural resources, and its scarcity is a prevalent issue in many parts of the world. Therefore, implementing smart water systems is not a luxury but a necessity. The primary role of smart water systems is to monitor, measure, and efficiently distribute water usage. Hauber-Davidson and Idris have designed a notable model in this field, the smart water metre [81]. These metres can detect water inflow and outflow and identify any potential leaks. Additionally, smart water metres can utilize data from smart river sensors and weather information to assist in flood prediction [4].
@@
Applications Related Works Applications Related Works
ECG Monitoring [82, 83] Glucose level monitoring [84, 85]
Temperature monitoring [86, 87] Blood pressure monitoring [88, 89]
Oxygen saturation monitoring [90, 91] Mood monitoring [92, 93]
Medication management [94, 95] Wheelchair management [96, 97]
Rehabilitation [98, 99] Fitness [100, 101]
Other notable application